Grade 11Physical Chemistry

Crude calcium carbide is made in an electric furnace by the following reaction:- CaO + 3C gives CaC2 + CO
The product contains 85 percent CaC2 and 15percent unreacted CaO.
Q1.How much pure CaO is to be added to the furnce charge for each 1000 kg of CaC2(pure) produced?
  1. How much CaO is to be added to the furnace charge for each 1000 kg of crude product?

Abhinav Acharya
12 Years agoGrade 11

1 Answer

Sumit Majumdar
12 Years ago
Dear student,
By equating the mole fraction for CaC2, we get:
\frac{\frac{1000\times 1000}{64.0994}}{\frac{1000\times 1000}{64.0994}+\frac{y}{56.0774}}=\frac{85}{100}
This gives that mass of CaO should be 154 kg.
Regards
Sumit